# Gaussian Process Implicit Surfaces as Control Barrier Functions for Safe Robot Navigation

### v1

@@ I INTRODUCTION
Safety-critical robotic systems are rapidly expanding across domains such as medicine, agriculture, warehouses, disaster response, and defense, where dependability is crucial since failures can endanger lives and cause major economic loss. Control barrier functions (CBFs), a level-set approach to enforcing safety [1], have been demonstrated on legged robots [2, 3], aerial vehicles [4], manipulators [5], and multi-agent systems [6]. CBFs guarantee forward invariance of a prescribed safe set using three elements: a candidate scalar function, a nominal system model, and a nominal control input. The candidate’s 0-superlevel set defines the CBF, from which quadratic program (QP) constraints rectify the nominal input into a safe control. However, most CBFs are hand-crafted from domain knowledge and heuristics, which is often infeasible in many real-world applications. This motivates our goal of achieving safe control for dynamical systems, particularly robots, by constructing CBFs through a data-driven approach.
@@
Data-driven approaches for constructing CBFs are actively pursued. Expert demonstrations of state and control trajectories have been used to generate CBFs [7], but relying on such demonstrations is impractical in unseen environments, and these methods lack hardware validation. Neural certificates have also been proposed [8, 9], providing formal correctness guarantees to learning-based controllers, yet they remain limited to offline training and simulation. Episodic learning has been applied to update controllers under safety constraints [10], but the focus was on modeling system uncertainty rather than constructing CBFs. Support vector machines have been used to characterize CBFs from sensor data [11], though they require carefully tuned weights and were demonstrated only in 2D simulation. Gaussian processes (GPs) have been employed in a LiDAR-based CBF formulation for 2D navigation of unicycle robots [12], and Bayesian meta-learning was studied for fast adaptation across environments in simulation [13]. In contrast to these works [12, 13], we characterize safe implicit surfaces in 3D using GPs, introduce sparse Gaussian CBFs to reduce complexity while retaining safety guarantees, and provide hardware results in 3D.
@@
To construct CBFs in 3D from data, we draw on signed distance functions (SDFs) as implicit surface representations. An SDF determines whether a point $\mathbf{x}\in\Omega$ lies inside or outside the set, and has been shown to be a natural representation for navigation and planning [14]. SDFs have also been used to construct CBFs with memory [15], though only with 2D sensing and in simulation. In contrast, we employ Gaussian processes (GPs) to form GP implicit surfaces (GPIS) [16], which have been studied for surface representation, navigation, shape estimation, and grasping [17, 18, 19, 20]. To the best of our knowledge, GPIS has not been explicitly used to synthesize CBFs. Motivated by SDFs for surface representation and CBFs for safe control, we propose a unified framework that directly infers implicit surfaces while enforcing safety control.
@@
The organization of the paper is: Section II presents preliminaries, Section III states the problem, and Section IV introduces our framework. Test cases are covered in Sections V and VI. Finally, conclusion is shared in Section VII.
@@ II BACKGROUND PRELIMINARIES
We model safety for a dynamical system using GPs as smooth surfaces, combining CBFs and GPIS to design probabilistic implicit surfaces. Below, we briefly review CBFs, GPs, and implicit surfaces (see [21], [22], [23] for details).
@@ II-A Control Barrier Function
Definition 1 (Control Barrier Function [21]).
@@ II-B Implicit Surfaces
The 0-level set $f_{\mathrm{is}}^{-1}(0)$ of an implicit function $f_{\mathrm{is}}$ defines a surface $\mathcal{S}\subset\mathbb{R}^{n}$, which in our case represents the safety boundary. Conversely, for any surface $\mathcal{S}$ in $\mathbb{R}^{n}$, there exists a function $f_{\mathrm{is}}:\mathbb{R}^{n}\to\mathbb{R}$ such that $f_{\mathrm{is}}^{-1}(0)=\mathcal{S}$ (Prop. 2 in [24]). Hence, given samples on $\mathcal{S}$, one can construct $f_{\mathrm{is}}$. We approximate the implicit surface of a volumetric object in $d=3$ ($d\leq n$) using a GP regressor, referred to in the literature as a Gaussian Process Implicit Surface (GPIS).
@@ II-C Gaussian Process Regression
where $\delta_{ij}=1$ if $i=j$ and 0 otherwise, $\mathbf{l}\in\mathbb{R}^{n}$ is the characteristic length scale, with $\mathbf{L}=\mathrm{diag}(\mathbf{l})\in\mathbb{R}^{n\times n}$. The signal scale and observation noise are $\sigma_{f}^{2}$ and $\sigma_{y}^{2}$ respectively. Together, these parameters constitute the SE kernel’s hyperparameters, $\Theta=\{\mathbf{L},\sigma_{f}^{2},\sigma_{y}^{2}\}$. These hyperparameters for a dataset can be optimized by maximizing the log marginal likelihood using quasi-Newton methods [22].
We are interested in constructing an implicit surface, which will later be our safety function of interest, for which we will carefully design scalar targets $y$ in Section IV-A. Given a set of $N$ data points, with input vectors $\mathbf{x}\in\mathbb{R}^{n}$, and scalar targets $y\in\mathbb{R}$, we compose the dataset $\mathcal{D}_{N}=\{\mathbf{X}_{N},\mathbf{y}_{N}\}$, where $\mathbf{X}_{N}=\{\mathbf{x}_{i}\}_{i=1}^{N}$ and $\mathbf{y}_{N}=\{y_{i}\}_{i=1}^{N}$. GPs can compute the posterior mean and variance for an arbitrary deterministic query point $\mathbf{x}_{*}\in\mathbb{R}^{n}$, by conditioning on previous measurements. The posterior mean $\mu\in\mathbb{R}$ and variance $\sigma^{2}\in\mathbb{R}$ are [22],
@@ III PROBLEM STATEMENT
This assumption is reasonable, since many sensing modalities (e.g., laser, haptic, or vision-based) provide point clouds and surface normals [19].
@@ IV PROPOSED METHODOLOGY
Here, we present our approach to approximate the GPIS $f_{\mathrm{is}}$ with a sparse Gaussian CBF $h_{\mathrm{sgp}}(\mathbf{x})$. GPs are preferable to models such as neural networks, polynomial chaos, or radial basis functions because their Bayesian nonparametric formulation provides flexible priors and closed-form probabilistic posterior estimates. Moreover, GPs naturally yield uncertainty through posterior variance, whereas neural networks require careful design to capture uncertainty [25].
@@ IV-B Sparse Gaussian Control Barrier Function
We use a GP prior on the desired safety function, $h_{\mathrm{sgp}}\sim\mathcal{GP}(0,k(\mathbf{x},\mathbf{x}^{\prime}))$ where the inputs and targets to the GP are $\mathbf{X}_{N}$ and $y_{N}$ as discussed above. By using the GP prior, we fully specify the candidate safety function and also present a sparse solution to our approach.
@@
Despite GPs being very powerful regressors, as the dataset grows larger, they become computationally intractable. GP prediction complexity has a cost of $\mathcal{O}(N^{3})$. Even if one stores the covariance matrix to save costs, the complexity per test case is $\mathcal{O}(N)$ for predictive mean and $\mathcal{O}(N^{2})$ for predictive variance. Hence, many sparse approximations of GPs have been developed to reduce the complexity cost while retaining accuracy [26, 27, 28]. We focus on the variant called Sparse Pseudo-Input Gaussian Process (SPGP) [27]. Here, we simply present its mean and variance:
@@
where $\hskip-4.26773pt\mathbf{k}_{M}(\mathbf{x}_{*})\hskip-8.5359pt=\hskip-8.5359pt\big[k(\overline{\mathbf{x}}_{1},\mathbf{x}_{*}),\ldots,k(\overline{\mathbf{x}}_{M},\mathbf{x}_{*})\big]^{\top}\hskip-14.22636pt\in\mathbb{R}^{M}$ is the covariance between $\overline{\mathbf{X}}_{M}$ and $\mathbf{x}_{*}$, $\mathbf{Q}_{M}=\mathbf{K}_{NM}^{\top}\big(\mathbf{\Lambda}_{N}+\sigma_{y}^{2}\mathbf{I}_{N}\big)^{-1}\mathbf{K}_{NM}+\mathbf{K}_{M}\in\mathbb{R}^{M\times M}$, $\mathbf{\Lambda}_{N}=\text{diag}\big[\mathbf{K}_{N}\ -\ \mathbf{K}_{NM}\ \mathbf{K}_{M}^{\ -1}\ \mathbf{K}_{MN}\big]\in\mathbb{R}^{N\times N}$ is a diagonal matrix, $\mathbf{K}_{M}\in\mathbb{R}^{M\times M}$ is the covariance between pairs of pseudo-inputs $\overline{\mathbf{X}}_{M}$, $\mathbf{P}_{M}=\mathbf{K}_{M}^{\ -1}-\mathbf{Q}_{M}^{\ -1}$, and $\big[\mathbf{K}_{NM}\big]_{(i,j)}=k(\mathbf{x}_{i},\overline{\mathbf{x}}_{j}),\ i\in\{1,...,N\},\ j\in\{1,...,M\}$ is the covariance matrix between $\mathbf{X}_{N}$ and $\overline{\mathbf{X}}_{M}$. Computation cost for $\mathbf{Q}_{M}$ is dominated by the inversion operation which is $\mathcal{O}(M^{2}N)$ [27]. By precomputing the inverse, the cost per test case is $\mathcal{O}(M)$ and $\mathcal{O}(M^{2})$ for predictive mean and variance respectively. We can jointly optimize for the kernel hyperparameters, $\Theta$, and pseudo-inputs, $\overline{\mathbf{X}}_{M}$, by maximizing the log marginal likelihood using quasi-Newton gradient methods:
@@
where $\overline{\mathbf{Q}}_{MN}=\mathbf{Q}_{M}^{-1}\ \mathbf{K}_{MN}\big(\mathbf{\Lambda}_{N}+\sigma_{y}^{2}\mathbf{I}_{N}\big)^{-1}\in\mathbb{R}^{M\times N}$. As can be seen from the formulation above, we fully realize the safety function using GPs, as well as exploit sparsity, to model any arbitrary volumetric object in an $n$-dimensional Euclidean space while addressing safety. The admissible control space for the sparse Gaussian CBF ensures that the system (1) remains forward invariant in the safe set characterized by $h_{\mathrm{sgp}}$ (Theorem 1 in [21]).
@@ IV-D Safe Control using Sparse Gaussian CBF
Given $\mathbf{u}_{\mathrm{nom}}$ and $\mathbf{x}_{\mathrm{des}}$ outside the safe set $\mathcal{S}$, the system would exit $\mathcal{S}$ and violate safety. To ensure forward invariance, $\mathbf{u}_{\mathrm{nom}}$ is rectified by solving an online QP with constraints from the Lie derivatives in (17) [21]:
@@ V-A1 Offline Training
The safety function is synthesized as the implicit surface of the bunny, with the Gaussian CBF formulation given in Section V-C. Training was performed with the gpml toolbox [29] on an Intel i7-9800X CPU (16 GB RAM, 4.4 $\mathrm{GHz}$). For the Gaussian CBF, 2,1782{,}178 surface points and normals were sampled, giving an average training time of $14.15$\mathrm{s}$\$. The sparse Gaussian CBF used one-fifth of these as pseudo-points, reducing training time to $9.46$\mathrm{s}$\$. Online training is described in Section V-D. Figure 2 shows the 0-isosurfaces: the Gaussian CBF captures fine facial contours, while the sparse variant preserves shape with less detail, showing the accuracy–speed trade-off.
@@ VI-A Experimental Hardware Overview
We use the Crazyflie 2.1, a $27$\mathrm{g}$\$ open-source quadrotor with a $15$\mathrm{g}$\$ payload limit [30]. Due to its limited payload limit, it is unsuitable for any onboard depth sensing; thus, we generate Gaussian CBF surfaces offline. State estimation runs onboard aided with the help of an external infrared positioning system [30], which requires line-of-sight. For collision-avoidance experiments we use an IKEA ADDE chair, chosen for its large backrest gap with multiple holes, enabling state estimation even when flying beneath or behind it (otherwise the lighthouse signal is occluded). GP training (with and without sparsity) and control rectification are executed on the same ground station as discussed in V-A.

### v2

@@ I INTRODUCTION
Safety-critical robotic systems are rapidly expanding across domains such as medicine, agriculture, warehouses, disaster response, and defense, where dependability is crucial since failures can endanger lives and cause major economic loss. Control barrier functions (CBFs), a level-set approach to enforcing safety [cbf_theory_Ames2019control], have been demonstrated on legged robots [cbf_legged_Grandia2021, cbf_legged_Tend2021], aerial vehicles [cbf_cascaded_Khan2019, cbf_racing_drones_Singletary2022], manipulators [cbf_manipulator_Rauscher2016], and multi-agent systems [cbf_swarm_consensus_Machida2021]. CBFs guarantee forward invariance of a prescribed safe set using three elements: a candidate scalar function, a nominal system model, and a nominal control input. The candidate’s 0-superlevel set defines the CBF, from which quadratic program (QP) constraints rectify the nominal input into a safe control. However, most CBFs are hand-crafted from domain knowledge and heuristics, which is often infeasible in many real-world applications. This motivates our goal of achieving safe control for dynamical systems, particularly robots, by constructing CBFs through a data-driven approach.
@@
Data-driven approaches for constructing CBFs are actively pursued. Expert demonstrations of state and control trajectories have been used to generate CBFs [cbf_robust_hybrid_Robey2021], but relying on such demonstrations is impractical in unseen environments, and these methods lack hardware validation. Neural certificates have also been proposed [neural_lyapunov_Gaby2021, neural_convex_Tsukamoto2020], providing formal correctness guarantees to learning-based controllers, yet they remain limited to offline training and simulation. Episodic learning has been applied to update controllers under safety constraints [cbf_episodic_biped_Csomay2021], but the focus was on modeling system uncertainty rather than constructing CBFs. Support vector machines have been used to characterize CBFs from sensor data [cbf_supervisedml_Srinivasan2020], though they require carefully tuned weights and were demonstrated only in 2D simulation. Gaussian processes (GPs) have been used to synthesize CBFs directly from data [cbf_gaussian_Khan2022]. Other studies involved using Bayesian meta-learning for fast adaptation across environments in simulation [cbf_bayesian_onboard_sensor_Hashimoto2023], and a LiDAR-based GP-CBF for 2D navigation of unicycle robots [cbf_gaussian_lidar_Keyumarsi2024]. In contrast to [cbf_gaussian_Khan2022, cbf_bayesian_onboard_sensor_Hashimoto2023, cbf_gaussian_lidar_Keyumarsi2024], we characterize safe implicit surfaces in 3D using GPs, introduce sparse Gaussian CBFs to reduce complexity while retaining safety guarantees, and provide hardware results in 3D.
@@
To construct CBFs in 3D from data, we draw on signed distance functions (SDFs) as implicit surface representations. An SDF determines whether a point $\mathbf{x}\in\Omega$ lies inside or outside the set, and has been shown to be a natural representation for navigation and planning [is_sdf_map_plan_Oleynikova2016]. SDFs have also been used to construct CBFs with memory [cbf_learning_sdf_Long2021], though only with 2D sensing and in simulation. In contrast, we employ Gaussian processes (GPs) to form GP implicit surfaces (GPIS) [is_gpis_Williams2006], which have been studied for surface representation, navigation, shape estimation, and grasping [is_gpis_esdf_Wu2021, is_gpis_mapping_Lee2019, is_gpis_shape_grasping_Dragiev2011, is_sparse_gpis_shape_tactile_Gandler2020]. To the best of our knowledge, GPIS has not been explicitly used to synthesize CBFs. Motivated by SDFs for surface representation and CBFs for safe control, we propose a unified framework that directly infers implicit surfaces while enforcing safety control.
@@ II BACKGROUND PRELIMINARIES
We model safety for a dynamical system using GPs as smooth surfaces, combining CBFs and GPIS to design probabilistic implicit surfaces. Below, we briefly review CBFs, GPs, and implicit surfaces (see [cbf_Ames2017], [gp_textbook_Rasmussen2003], [is_textbook_levelset_Osher2004] for details).
@@ II-A Control Barrier Function
Definition 1 (Control Barrier Function [cbf_Ames2017]).
@@ II-B Implicit Surfaces
The 0-level set $f_{\mathrm{is}}^{-1}(0)$ of an implicit function $f_{\mathrm{is}}$ defines a surface $\mathcal{S}\subset\mathbb{R}^{n}$, which in our case represents the safety boundary. Conversely, for any surface $\mathcal{S}$ in $\mathbb{R}^{n}$, there exists a function $f_{\mathrm{is}}:\mathbb{R}^{n}\to\mathbb{R}$ such that $f_{\mathrm{is}}^{-1}(0)=\mathcal{S}$ (Prop. 2 in [is_textbook_Do2016]). Hence, given samples on $\mathcal{S}$, one can construct $f_{\mathrm{is}}$. We approximate the implicit surface of a volumetric object in $d=3$ ($d\leq n$) using a GP regressor, referred to in the literature as a Gaussian Process Implicit Surface (GPIS).
@@ II-C Gaussian Process Regression
where $\delta_{ij}=1$ if $i=j$ and 0 otherwise, $\mathbf{l}\in\mathbb{R}^{n}$ is the characteristic length scale, with $\mathbf{L}=\mathrm{diag}(\mathbf{l})\in\mathbb{R}^{n\times n}$. The signal scale and observation noise are $\sigma_{f}^{2}$ and $\sigma_{y}^{2}$ respectively. Together, these parameters constitute the SE kernel’s hyperparameters, $\Theta=\{\mathbf{L},\sigma_{f}^{2},\sigma_{y}^{2}\}$. These hyperparameters for a dataset can be optimized by maximizing the log marginal likelihood using quasi-Newton methods [gp_textbook_Rasmussen2003].
We are interested in constructing an implicit surface, which will later be our safety function of interest, for which we will carefully design scalar targets $y$ in Section IV-A. Given a set of $N$ data points, with input vectors $\mathbf{x}\in\mathbb{R}^{n}$, and scalar targets $y\in\mathbb{R}$, we compose the dataset $\mathcal{D}_{N}=\{\mathbf{X}_{N},\mathbf{y}_{N}\}$, where $\mathbf{X}_{N}=\{\mathbf{x}_{i}\}_{i=1}^{N}$ and $\mathbf{y}_{N}=\{y_{i}\}_{i=1}^{N}$. GPs can compute the posterior mean and variance for an arbitrary deterministic query point $\mathbf{x}_{*}\in\mathbb{R}^{n}$, by conditioning on previous measurements. The posterior mean $\mu\in\mathbb{R}$ and variance $\sigma^{2}\in\mathbb{R}$ are [gp_textbook_Rasmussen2003],
@@ III PROBLEM STATEMENT
This assumption is reasonable, since many sensing modalities (e.g., laser, haptic, or vision-based) provide point clouds and surface normals [is_gpis_shape_grasping_Dragiev2011].
@@ IV PROPOSED METHODOLOGY
Here, we present our approach to approximate the GPIS $f_{\mathrm{is}}$ with a sparse Gaussian CBF $h_{\mathrm{sgp}}(\mathbf{x})$. GPs are preferable to models such as neural networks, polynomial chaos, or radial basis functions because their Bayesian nonparametric formulation provides flexible priors and closed-form probabilistic posterior estimates. Moreover, GPs naturally yield uncertainty through posterior variance, whereas neural networks require careful design to capture uncertainty [nn_survey_Gawlikowski2021].
@@ IV-B Sparse Gaussian Control Barrier Function
We use a GP prior on the desired safety function, $h_{\mathrm{sgp}}\sim\mathcal{GP}(0,k(\mathbf{x},\mathbf{x}^{\prime}))$ where the inputs and targets to the GP are $\mathbf{X}_{N}$ and $y_{N}$ as discussed above. By using the GP prior, we fully specify the candidate safety function, unlike [cbf_gaussian_variance_Khan2021] where only the posterior variance was modeled by GP. Additionally, we also present a sparse approximation to our approach.
@@
Despite GPs being very powerful regressors, as the dataset grows larger, they become computationally intractable. GP prediction complexity has a cost of $\mathcal{O}(N^{3})$. Even if one stores the covariance matrix to save costs, the complexity per test case is $\mathcal{O}(N)$ for predictive mean and $\mathcal{O}(N^{2})$ for predictive variance. Hence, many sparse approximations of GPs have been developed to reduce the complexity cost while retaining accuracy [gp_power_ep_Minka2001, gp_sparse_fitc_Snelson2006, gp_sparse_vfe_Titsias2009]. We focus on the variant called Sparse Pseudo-Input Gaussian Process (SPGP) [gp_sparse_fitc_Snelson2006]. Here, we simply present its mean and variance:
@@
where $\hskip-4.26773pt\mathbf{k}_{M}(\mathbf{x}_{*})\hskip-8.5359pt=\hskip-8.5359pt\big[k(\overline{\mathbf{x}}_{1},\mathbf{x}_{*}),\ldots,k(\overline{\mathbf{x}}_{M},\mathbf{x}_{*})\big]^{\top}\hskip-14.22636pt\in\mathbb{R}^{M}$ is the covariance between $\overline{\mathbf{X}}_{M}$ and $\mathbf{x}_{*}$, $\mathbf{Q}_{M}=\mathbf{K}_{NM}^{\top}\big(\mathbf{\Lambda}_{N}+\sigma_{y}^{2}\mathbf{I}_{N}\big)^{-1}\mathbf{K}_{NM}+\mathbf{K}_{M}\in\mathbb{R}^{M\times M}$, $\mathbf{\Lambda}_{N}=\text{diag}\big[\mathbf{K}_{N}\ -\ \mathbf{K}_{NM}\ \mathbf{K}_{M}^{\ -1}\ \mathbf{K}_{MN}\big]\in\mathbb{R}^{N\times N}$ is a diagonal matrix, $\mathbf{K}_{M}\in\mathbb{R}^{M\times M}$ is the covariance between pairs of pseudo-inputs $\overline{\mathbf{X}}_{M}$, $\mathbf{P}_{M}=\mathbf{K}_{M}^{\ -1}-\mathbf{Q}_{M}^{\ -1}$, and $\big[\mathbf{K}_{NM}\big]_{(i,j)}=k(\mathbf{x}_{i},\overline{\mathbf{x}}_{j}),\ i\in\{1,...,N\},\ j\in\{1,...,M\}$ is the covariance matrix between $\mathbf{X}_{N}$ and $\overline{\mathbf{X}}_{M}$. Computation cost for $\mathbf{Q}_{M}$ is dominated by the inversion operation which is $\mathcal{O}(M^{2}N)$ [gp_sparse_fitc_Snelson2006]. By precomputing the inverse, the cost per test case is $\mathcal{O}(M)$ and $\mathcal{O}(M^{2})$ for predictive mean and variance respectively. We can jointly optimize for the kernel hyperparameters, $\Theta$, and pseudo-inputs, $\overline{\mathbf{X}}_{M}$, by maximizing the log marginal likelihood using quasi-Newton gradient methods:
@@
where $\overline{\mathbf{Q}}_{MN}=\mathbf{Q}_{M}^{-1}\ \mathbf{K}_{MN}\big(\mathbf{\Lambda}_{N}+\sigma_{y}^{2}\mathbf{I}_{N}\big)^{-1}\in\mathbb{R}^{M\times N}$. As can be seen from the formulation above, we fully realize the safety function using GPs, as well as exploit sparsity, to model any arbitrary volumetric object in an $n$-dimensional Euclidean space while addressing safety. The admissible control space for the sparse Gaussian CBF ensures that the system (1) remains forward invariant in the safe set characterized by $h_{\mathrm{sgp}}$ (Theorem 1 in [cbf_Ames2017]).
@@ IV-D Safe Control using Sparse Gaussian CBF
Given $\mathbf{u}_{\mathrm{nom}}$ and $\mathbf{x}_{\mathrm{des}}$ outside the safe set $\mathcal{S}$, the system would exit $\mathcal{S}$ and violate safety. To ensure forward invariance, $\mathbf{u}_{\mathrm{nom}}$ is rectified by solving an online QP with constraints from the Lie derivatives in (17) [cbf_Ames2017]:
@@ V-A1 Offline Training
The safety function is synthesized as the implicit surface of the bunny, with the Gaussian CBF formulation given in Section V-C. Training was performed with the gpml toolbox [gp_toolbox_Rasmussen2010] on an Intel i7-9800X CPU (16 GB RAM, 4.4 $\mathrm{GHz}$). For the Gaussian CBF, 2,1782{,}178 surface points and normals were sampled, giving an average training time of $14.15$\mathrm{s}$\$. The sparse Gaussian CBF used one-fifth of these as pseudo-points, reducing training time to $9.46$\mathrm{s}$\$. Online training is described in Section V-D. Figure 2 shows the 0-isosurfaces: the Gaussian CBF captures fine facial contours, while the sparse variant preserves shape with less detail, showing the accuracy–speed trade-off.
@@ VI-A Experimental Hardware Overview
We use the Crazyflie 2.1, a $27$\mathrm{g}$\$ open-source quadrotor with a $15$\mathrm{g}$\$ payload limit [crazyflie]. Due to its limited payload limit, it is unsuitable for any onboard depth sensing; thus, we generate Gaussian CBF surfaces offline. State estimation runs onboard aided with the help of an external infrared positioning system [crazyflie], which requires line-of-sight. For collision-avoidance experiments we use an IKEA ADDE chair, chosen for its large backrest gap with multiple holes, enabling state estimation even when flying beneath or behind it (otherwise the lighthouse signal is occluded). GP training (with and without sparsity) and control rectification are executed on the same ground station as discussed in V-A.